it takes Ellen 9 hours to drive to a friend's cabin. Her sister Barb can make the drive in 8 hours by traveling 7mph faster than Ellen. What is Barb's average speed?

Answer by mananth(16946)


Let the distance be y
it is the same for both
the speeds are different
y/x=9
y=9x
y/x+7 =8
y=8(x+7)
9x=8(x+7)
9x=8x+56
9x-8x=56
x=56 Ellen's speed
Barb's speed = 56+7 = 63 mph